A hiker begins a trip by first walking 30 km southeast from her car. She stops and sets up her tent for the night. On the second day, she walks 40 km in a direction 70° north of east, at which point she discovers a forest ranger’s tower.



1
Expert's answer
2021-02-12T06:11:47-0500

a) According to the directions distribution below


the first displacement vector is given as follows:


"\mathbf{d}_1 = 30(\cos45\degree, -\cos 45\degree) = (15,-15)"

The second displacement is given as follows:


"\mathbf{d}_2 = (40\cos70\degree, 40\sin70\degree)"

b) The resultant displacement is:


"\mathbf{R} = \mathbf{d}_1 + \mathbf{d}_2 = (15+40\cos 70\degree,-15+40\sin 70\degree)\\
\mathbf{R}\approx (28.68,22.59)"

c) The magnitude is:


"R = \sqrt{28.68^2+22.59^2} \approx 36.51m"

The direction is:


"\theta = \arctan\left( \dfrac{22.59}{28.68} \right) \approx 38\degree"

which is "38\degree" north of east.


Answer. a) "(15,-15)", "(40\cos70\degree, 40\sin70\degree)", b) "(28.68,22.59)", c) "36.51" m, "38\degree" north of east.